# RobEthiChor: Automated Context-aware Ethics-based Negotiation for Autonomous Robots

### v1

@@ 1 Introduction
The presence of autonomous decision-making systems is growing at a fast pace and is poised to have a critical impact on society [1, 2]. Increasingly, these systems will be acting on behalf of humans [3] and making decisions autonomously without the need for human intervention [4]. As of today, however, human values and ethics are mostly not considered by the digital systems that are making decisions for them, leading to a lack of trust [5, 6, 7]. In particular, the autonomous systems that we have today are not able to consider the ethical preferences of end users, which can indeed be used to customize their decision-making process. To truly serve our needs and become an integral part of our lives, these systems must expand their abilities to integrate mechanisms that enable ethical considerations in their decision-making process [8, 9, 10, 11, 12]. This calls for enhancing the decision-making to consider the personal ethical preferences of users and take into account their morals and beliefs, which may vary over contexts and are unique for each individual in society [12, 13, 14, 15, 16, 17]. For instance, in fields such as healthcare, where service robots help in patient care, smart home systems, where robots provide personal assistance, autonomous vehicles, etc., it is essential for these systems to take into account the ethical preferences of their users to make ethically-informed decisions.
@@
Although several studies [18, 19, 20, 21, 22, 23] consider ethics in their decision-making, these systems adhere to ethical principles established by system designers (as detailed below in Section 11). Hence, due to the absence of a universal ethic, its intrinsic subjectivity and capillarity [24], the challenge of “personalizing” autonomous systems with the moral beliefs of end users (e.g., through an ethical profile) for the purpose of customizing their decision-making remains mostly unaddressed [25]. For instance, how should the system make a decision that accounts for the user’s benevolence towards the elderly or people with health problems? In this direction, the work in [26] delves into the foundational principles of addressing user ethical concerns in autonomous systems: individuals need to be able to exercise (some degree of) control over the decisions that autonomous systems make on their behalf, i.e., autonomous systems, in their behavior, need to consider user’s moral beliefs. Empowering the user with personalized software connectors that are able to reflect the user’s moral values in the interactions with digital systems is also the approach taken in the Exosoul project [27]. The approach has been applied to privacy profiles [28, 29] and ethical profiles [30]. Following this line of work, in [31], the authors emphasize the need for user empowerment within the framework of autonomous systems, which demand mechanisms for balancing diverse needs, values, and ethics at the individual, community, and societal levels.
@@
To address these challenges, this paper presents RobEthiChor, an approach that makes autonomous systems capable of exploiting user ethical preferences to regulate their autonomous behavior and support decision-making through ethics-based negotiation. Specifically, we target autonomous systems that act on behalf of humans [3] and, building on the principles of Utilitarianism [52, 53, 54], exploit a user’s ethical profile and a context model to personalize their autonomy and, upon a resource contention, negotiate to reach an agreement. The approach features a domain-agnostic reference architecture as a template solution for realizing autonomous systems capable of leveraging the ethical profiles of end users and their contextual status during negotiation, so to ensure that the decision is aligned with the preferences of all parties involved. In this paper, we also present RobEthiChor-Ros, a specific instance of our reference architecture implementing the proposed automated context-aware ethics-based negotiation on top of the Robot Operating System (ROS). The full implementation of RobEthiChor-Ros is available [^1] https://github.com/gianlucafilippone/robethichor-ros, and the experimental results show that the approach effectively accomplishes the negotiation and allows systems to behave according to user ethical preferences. The replication package permits reproducing the experiments [^2] https://github.com/RoboChor/robethichor-ethics-based-negotiation. In summary, main contributions are:
@@
The paper is organized as follows. Section 2 introduces the main ingredients of our approach. Section 3 presents a motivating scenario, and Section 4 details the proposed solution. Section 5 describes the reference architecture, and Section 6 presents its implementation. Section 7 presents the evaluation of the approach and its results. Sections 8 and 10 discuss relevant aspects of our approach and highlight limitations and threats to validity. Section 11 discusses related work, and Section 12 concludes the paper with future steps.
@@ 2.1 Digital ethics
We employ the notion of digital ethics, as proposed by Floridi [55], which differentiates between hard ethics, referring to the ethical rules described by the higher authorities which are, mostly often, domain-specific and, in principle, commonly accepted, and soft ethics, which encompasses the individual’s ethical preferences [55], being individuals single people, groups, associations, etc. We consider that autonomous systems comply with the hard ethics of the domain/context and negotiate based on their user’s soft ethical preferences to (possibly) reach an ethical agreement that aligns with the ethical preferences of all parties involved. The user’s soft ethical preferences that we consider in this paper are inspired by the notion of dispositions [56]. In its most general philosophical acceptation, dispositions refer to the capacity of an object to behave differently under different circumstances [57]. For example, water exhibits different behaviors according to environmental conditions, i.e., it becomes solid at freezing temperatures, liquid under moderate temperatures, and vaporizes upon heating, adapting its form naturally to the environment. Similarly, a given user may act differently in a hospital compared to an airport based on his/her contextual conditions/status/dispositions (as we will describe in Section 3). Dispositions are gradable properties, associated with the context, which can be instantiated by any kind of entity [58]. Grades are assigned to dispositional properties according to the importance that the users attribute to each disposition, i.e., the willingness of a user to behave in a certain way under specific circumstances, thereby reflecting their preferences/values within a specific context. More recently, the problem of how to embed human values in the design and implementation of software systems, notably the problem of how to operationalize human values, has received increasing attention in the software engineering community [27, 59, 60, 61]. Beyond our approach to modeling ethical preferences through dispositions, all those approaches can be well exploited to provide user ethical profiles that enjoy similar properties.
@@ 3.2 Scenario description
A visual representation of the scenario is presented in Figure 1. RobAssist $A$ and RobAssist $B$ are programmed with the mission of transporting passengers to their gates using service wheelchairs (Figure 1a). $Alice$ is an athlete who suffered a leg injury and hence would prefer assistance to move through the airport. $Bob$ is an elderly traveler and is running late for the flight, which is boarding shortly. $Bob$ requires assistance in rapidly moving to the departure gate. The two assistive robots RobAssist $A$ and RobAssist $B$ are assigned respectively to $Alice$ and $Bob$ to provide them the required assistance. When encountering RobAssist $A$ and RobAssist $B$, $Alice$ and $Bob$ use their mobile phones to scan the robots and provide their digital ethical profiles through the app RobEthiApp. Additionally, they use the app to inform the assistive robots of their current status by selecting a set of possible conditions that apply to them, e.g., elderly, injured, etc. (Figure 1b). [^3] Note that, using the companion app, profiles can be defined beforehand and reused for different scenarios and any compatible system. Similarly, user status conditions can be provided to the app beforehand and updated on the fly, if desired. $Alice$ and $Bob$ further provide their flight details by scanning the boarding cards, hence allowing the robots to dynamically identify other relevant contextual information, such as flight status updates, etc (Figure 1c).
@@ 4.1 Context model
[…]
@@ 4.2 User status model
[…]
@@
$elderly$ $d_{1}$ Give precedence to elderly people 2 2
$injured$ $d_{2}$ Give precedence to injured people 3 4
$boarding\_emergency$ $d_{3}$ Give precedence to the ones in emergency 5 5
$pet\_owner$ $d_{4}$ Give precedence to the ones traveling with pet 2 N/A
$reduced\_oxygen$ $d_{5}$ Give precedence to the ones with serious medical conditions N/A 3
@@ 4.3 Ethics profile(s)
For example, given $\mathcal{D}$=$=\{d_{1},d_{2},d_{3},d_{4},d_{5}\}$ in Table 1, and the airport and hospital contexts $\mathcal{C}^{A}_{a}$ and $\mathcal{C}^{A}_{h}$ above, two possible context-related profiles for $Alice$ are $E^{A}_{\pi}(\mathcal{C}^{A}_{a})$$=$$(\mathcal{D},$$\preceq_{\mathcal{C}^{A}_{a}}$))) and $E^{A}_{\pi}(\mathcal{C}^{A}_{h})$$=$$(\mathcal{D},$ $\preceq_{\mathcal{C}^{A}_{h}}$))). According to the assigned grades [^6] Importantly, the given order of grades is purely coincidental and only for the purpose of our motivating example., we have $d_{1}$$\preceq_{\mathcal{C}^{A}_{a}}$$d_{4}$$\preceq_{\mathcal{C}^{A}_{a}}$$d_{2}$ $\preceq_{\mathcal{C}^{A}_{a}}$$d_{3}$, and $d_{1}$$\preceq_{\mathcal{C}^{A}_{h}}$$d_{5}$$\preceq_{\mathcal{C}^{A}_{h}}$$d_{2}$$\preceq_{\mathcal{C}^{A}_{h}}$$d_{3}$. Note that, it can be the case that, in a given context, only a subset of dispositions applies.
@@ 4.4 Task ethical implication
Back to our example, within the airport context $\mathcal{C}^{A}_{a}$, the task $t_{4}$ = take elevator has an impact on the disposition $d_{2}$ = give precedence to injured people (which has been aptly activated by the $Alice$’s injured status in $\mathcal{S}_{a}^{A}$): hence, for $Alice$, $E^{A}_{\pi}(t_{4},\mathcal{C}^{A}_{a},\mathcal{S}_{a}^{A})$$=$$\{d_{2}\}$. Similarly, based on $Bob$’s elderly and emergency status $\mathcal{S}_{a}^{B}$, the same task $t_{4}$ has an impact on the dispositions $d_{1}$ = give precedence to elderly people and $d_{3}$ = give precedence to the ones in emergency; hence, for $Bob$, $E^{B}_{\pi}(t_{4},\mathcal{C}_{a},$ $\mathcal{S}_{a}^{B})=\{d_{1},d_{3}\}$. Note that, thanks to the current user status, we can distinguish if an ethical implication holds in a particular negotiation case.
@@ 4.5 Task ethical impact
Given the ethical implication $E_{\pi}(t,\mathcal{C},\mathcal{S})$, the ethical impact of the task $t\in\mathcal{T}$ in the context $\mathcal{C}$, given the user status $\mathcal{S}$, is defined as $\mathcal{T}_{\varepsilon\iota}(t,\mathcal{C},\mathcal{S})=\sum_{d\in E_{\pi}(t,\mathcal{C},\mathcal{S})}\lfloor t,\mathcal{C},d\rceil$, where $\lfloor t,\mathcal{C},d\rceil=rank(d,\mathcal{C})$ if $d$ manifests with a positive impact when performing $t$ in the context $\mathcal{C}$, $-rank(d,\mathcal{C})$ if $d$ manifests with a negative impact, $N/A$ if $d$ does not manifest with the task $t$. Still referring to Table 1, for $Alice$, we have $\mathcal{T}_{\varepsilon\iota}(t_{4},\mathcal{C}^{A}_{a},\mathcal{S}_{a}^{A})=3$; whereas, for $Bob$, $\mathcal{T}_{\varepsilon\iota}(t_{4},\mathcal{C}_{a},\mathcal{S}_{a}^{B})=7$
@@ 4.7 Utility function
For our scenario, upon the first negotiation iteration, $Bob$ is the sender, $Alice$ the receiver, and the first offer $o_{1}^{B\rightarrow A}=\langle\{t_{4}\},\{\langle elderly,$ $true\rangle\}\rangle$ is compared with the $Alice$’s $\bar{o}=\langle\{t_{4}\},$$\{\langle injured,true\rangle,$ $\langle flight\_anxiety,true\rangle\}\rangle$ (see the $Alice$’s profile in the context $\mathcal{C}^{A}_{a}$ in Table 1). As per the grades in the table, the utility function has value $\mathcal{U}(o_{1},\bar{o})$ $=$ $\mathcal{O}_{\varepsilon\iota}(o_{1},\mathcal{C}^{A}_{a})-\mathcal{O}_{\varepsilon\iota}(\bar{o},\mathcal{C}^{A}_{a})$ $=$ $-2$ 4 = - 2. From the point of view of $Alice$, by considering the only status conditions disclosed by $Bob$ at the first iteration, a negative utility means that RobAssist $B$ performing the offered task $t_{4}$ for $Bob$ is less ethical than RobAssist $A$ performing $t_{4}$ for $Alice$.

### v2

@@ 1 Introduction
The presence of autonomous decision-making systems is growing at a fast pace and is poised to have a critical impact on society [1, 2]. Increasingly, these systems will be acting on behalf of humans [3] and making decisions autonomously without the need for human intervention [4]. As of today, however, human values and ethics are mostly not considered by the digital systems that are making decisions for them, leading to a lack of trust, which in fact requires considering not only technical and legal aspects, but also ethical ones [5, 6, 7, 8]. In particular, the autonomous systems that we have today are not able to consider the ethical preferences of end users, which can indeed be used to customize their decision-making process. This inability undermines their ethical trustworthiness, which, alongside technical trustworthiness, forms a crucial component of an overall trustworthy system [9, 10]. To truly serve our needs and become an integral part of our lives, these systems must expand their abilities to integrate mechanisms that enable ethical considerations in their decision-making process [11, 12, 13, 14, 15]. This calls for enhancing the decision-making to consider the personal ethical preferences of users and take into account their morals and beliefs, which may vary over contexts and are unique for each individual in society [15, 16, 17, 18, 19, 20]. For instance, in fields such as healthcare, where service robots help in patient care, smart home systems, where robots provide personal assistance, autonomous vehicles, etc., it is essential for these systems to take into account the ethical preferences of their users to make ethically-informed decisions. Engineering autonomous systems that can reflect on ethical considerations will not only enable their integration into society but also foster trust toward these systems [9, 21].
@@
Although several studies [22, 23, 24, 25, 26, 27] consider ethics in their decision-making, these systems adhere to ethical principles established by system designers (as detailed below in Section 11). Hence, due to the absence of a universal ethic, its intrinsic subjectivity and capillarity [28], the challenge of “personalizing” autonomous systems with the moral beliefs of end users (e.g., through an ethical profile) for the purpose of customizing their decision-making remains mostly unaddressed [10]. For instance, how should the system make a decision that accounts for the user’s benevolence towards the elderly or people with health problems?
In this direction, the work in [29] delves into the foundational principles of addressing user ethical concerns in autonomous systems: individuals need to be able to exercise (some degree of) control over the decisions that autonomous systems make on their behalf, i.e., autonomous systems, in their behavior, need to consider user’s moral beliefs. Empowering the user with personalized software connectors that are able to reflect the user’s moral values in the interactions with digital systems is also the approach taken in the Exosoul project [30]. The approach has been applied to privacy profiles [31, 32] and ethical profiles [33]. Following this line of work, in [34], the authors emphasize the need for user empowerment within the framework of autonomous systems, which demand mechanisms for balancing diverse needs, values, and ethics at the individual, community, and societal levels.
@@
To address these challenges, this paper presents RobEthiChor, an approach that makes autonomous systems capable of exploiting user ethical preferences to regulate their autonomous behavior and support decision-making through ethics-based negotiation. The RobEthiChor approach aims to contribute towards the design and development of trustworthy autonomous systems capable of replicating the behavior of humans while making ethically informed decisions on their behalf. The goal is to promote the system’s trustworthiness by adopting an architecture that, supporting automated ethical negotiation, allows systems to take into account ethical considerations at runtime [55, 56, 57]. Specifically, we target autonomous systems that act on behalf of humans [3] and, building on the principles of Utilitarianism [58, 59, 60], exploit a user’s ethical profile and a context model to personalize their autonomy and, upon a resource contention, negotiate to reach an agreement. The approach features a domain-agnostic reference architecture as a template solution for realizing autonomous systems capable of leveraging the ethical profiles of end users and their contextual status during negotiation, so to ensure that the decision is aligned with the preferences of all parties involved. In this paper, we also present RobEthiChor-Ros, a specific instance of our reference architecture implementing the proposed automated context-aware ethics-based negotiation on top of the Robot Operating System (ROS). The full implementation of RobEthiChor-Ros is available [^1] https://github.com/gianlucafilippone/robethichor-ros, and the experimental results show that the approach effectively accomplishes the negotiation and allows systems to behave according to user ethical preferences. The replication package permits reproducing the experiments [^2] https://doi.org/10.5281/zenodo.17449923. In summary, main contributions are:
@@
The rest of the paper is organized as follows. Section 2 introduces the main ingredients of our approach. Section 3 presents a motivating scenario, and Section 4 details the proposed solution. Section 5 describes the reference architecture, and Section 6 presents its implementation. Section 7 presents the evaluation of the approach and its results. Sections 8 and 10 discuss relevant aspects of our approach and highlight limitations and threats to validity. Section 11 discusses related work, and Section 12 concludes the paper with future steps.
@@ 2.1 Digital ethics
We employ the notion of digital ethics, as proposed by Floridi [61], which differentiates between hard ethics, referring to the ethical rules described by the higher authorities which are, mostly often, domain-specific and, in principle, commonly accepted, and soft ethics, which encompasses the individual’s ethical preferences [61], being individuals single people, groups, associations, etc. We consider that autonomous systems comply with the hard ethics of the domain/context and negotiate based on their users’ soft ethical preferences to (possibly) reach an ethical agreement that aligns with the ethical preferences of all parties involved. The user’s soft ethical preferences that we consider in this paper are inspired by the notion of dispositions [62]. In its most general philosophical acceptation, dispositions refer to the capacity of an object to behave differently under different circumstances [63]. For example, water exhibits different behaviors according to environmental conditions, i.e., it becomes solid at freezing temperatures, liquid under moderate temperatures, and vaporizes upon heating, adapting its form naturally to the environment. Similarly, a given user may act differently in a hospital compared to an airport based on his/her contextual conditions/status/dispositions (as we will describe in Section 3). Dispositions are gradable properties, associated with the context, which can be instantiated by any kind of entity [64]. Grades are assigned to dispositional properties according to the importance that the users attribute to each disposition, i.e., the willingness of a user to behave in a certain way under specific circumstances, thereby reflecting their preferences/values within a specific context. More recently, the problem of how to embed human values in the design and implementation of software systems, notably the problem of how to operationalize human values, has received increasing attention in the software engineering community [30, 65, 66, 57]. Beyond our approach to modeling ethical preferences through dispositions, all those approaches can be well exploited to provide user ethical profiles that enjoy similar properties.
@@ 3.2 Scenario description
A visual representation of the scenario is presented in Figure 1. RobAssist $A$ and RobAssist $B$ are programmed with the mission of transporting passengers to their gates using service wheelchairs (Figure 1a). $Alice$ is an athlete who suffered a leg injury and hence would prefer assistance to move through the airport. $Bob$ is an elderly traveler and is running late for the flight, which is boarding shortly. $Bob$ requires assistance in rapidly moving to the departure gate. The two assistive robots RobAssist $A$ and RobAssist $B$ are assigned respectively to $Alice$ and $Bob$ to provide them the required assistance. When encountering RobAssist $A$ and RobAssist $B$, AliceAlice and $Bob$ use their mobile phones to scan the robots and provide their digital ethical profiles through the app RobEthiApp. Additionally, they use the app to inform the assistive robots of their current status by selecting a set of possible conditions that apply to them, e.g., elderly, injured, etc. (Figure 1b). [^3] Note that, using the companion app, profiles can be defined beforehand and reused for different scenarios and any compatible system. Similarly, user status conditions can be provided to the app beforehand and updated on the fly, if desired. $Alice$ and $Bob$ further provide their flight details by scanning the boarding cards, hence allowing the robots to dynamically identify other relevant contextual information, such as flight status updates, etc (Figure 1c).
@@ 4.1 Context model
A context is a set of pairs $\mathcal{C}=\{\langle a_{1},v_{1}\rangle,\cdots,\langle a_{n},$ $v_{n}\rangle\}$, nn>>0, where each $\langle a_{i},v_{i}\rangle$, $1\leqii\leqnn$, is a pair associating the context attribute name $a_{i}$ with its value $v_{i}$. Attribute names are keywords that, by default, are associated with the type of the attribute, e.g., boolean, string, integer, address/coordinates, enumerated.
@@ 4.2 User status model
Similarly to the context model above, user status is modeled as a set of pairs $\mathcal{S}=\{\langle c_{1},v_{1}\rangle,\cdots,\langle c_{m},v_{m}\rangle\}$, mm>>0, where each $\langle c_{j},v_{j}\rangle$, $1\leqjj\leqmm$, is a pair associating the name of the user condition $c_{j}$ with its value $v_{j}\in$$\{true,false\}$. Thus, differently from the context model, the type of the pairs is always boolean.
@@
elderlyelderly $d_{1}$ Give precedence to elderly people 2 2
injuredinjured $d_{2}$ Give precedence to injured people 3 4
boarding_emergencyboarding\_emergency $d_{3}$ Give precedence to the ones in emergency 5 5
pet_ownerpet\_owner $d_{4}$ Give precedence to the ones traveling with pet 2 N/A
reduced_oxygenreduced\_oxygen $d_{5}$ Give precedence to the ones with serious medical conditions N/A 3
@@ 4.3 Ethics profile(s)
For example, given $\mathcal{D}$$\{d_{1},d_{2},d_{3},d_{4},d_{5}\}$ in Table 1, and the airport and hospital contexts $\mathcal{C}^{A}_{a}$ and $\mathcal{C}^{A}_{h}$ above, two possible context-related profiles for $Alice$ are $E^{A}_{\pi}(\mathcal{C}^{A}_{a})$$(\mathcal{D},$$\preceq_{\mathcal{C}^{A}_{a}}$)) and $E^{A}_{\pi}(\mathcal{C}^{A}_{h})$$(\mathcal{D},$ $\preceq_{\mathcal{C}^{A}_{h}}$)). According to the assigned grades [^6] Importantly, the given order of grades is purely coincidental and only for the purpose of our motivating example., we have $d_{1}$$\preceq_{\mathcal{C}^{A}_{a}}$$d_{4}$$\preceq_{\mathcal{C}^{A}_{a}}$$d_{2}$ $\preceq_{\mathcal{C}^{A}_{a}}$$d_{3}$, and $d_{1}$$\preceq_{\mathcal{C}^{A}_{h}}$$d_{5}$$\preceq_{\mathcal{C}^{A}_{h}}$$d_{2}$$\preceq_{\mathcal{C}^{A}_{h}}$$d_{3}$. Note that, it can be the case that, in a given context, only a subset of dispositions applies.
@@ 4.4 Task ethical implication
Back to our example, within the airport context $\mathcal{C}^{A}_{a}$, the task $t_{4}$ = take elevator has an impact on the disposition $d_{2}$ = give precedence to injured people (which has been aptly activated by the $Alice$’s injured status in $\mathcal{S}_{a}^{A}$): hence, for $Alice$, $E^{A}_{\pi}(t_{4},\mathcal{C}^{A}_{a},\mathcal{S}_{a}^{A})$$\{d_{2}\}$. Similarly, based on $Bob$’s elderly and emergency status $\mathcal{S}_{a}^{B}$, the same task $t_{4}$ has an impact on the dispositions $d_{1}$ = give precedence to elderly people and $d_{3}$ = give precedence to the ones in emergency; hence, for $Bob$, $E^{B}_{\pi}(t_{4},\mathcal{C}_{a},$ $\mathcal{S}_{a}^{B})=\{d_{1},d_{3}\}$. Note that, thanks to the current user status, we can distinguish if an ethical implication holds in a particular negotiation case.
@@ 4.5 Task ethical impact
Given the ethical implication $E_{\pi}(t,\mathcal{C},\mathcal{S})$, the ethical impact of the task $t\in\mathcal{T}$ in the context $\mathcal{C}$, given the user status $\mathcal{S}$, is defined as $\mathcal{T}_{\varepsilon\iota}(t,\mathcal{C},\mathcal{S})=\sum_{d\in E_{\pi}(t,\mathcal{C},\mathcal{S})}\lfloor t,\mathcal{C},d\rceil$, where $\lfloor t,\mathcal{C},d\rceil=rank(d,\mathcal{C})$ if $d$ manifests with a positive impact when performing $t$ in the context $\mathcal{C}$, $rank(d,\mathcal{C})$ if $d$ manifests with a negative impact, $N/A$ if $d$ does not manifest with the task $t$. Still referring to Table 1, for $Alice$, we have $\mathcal{T}_{\varepsilon\iota}(t_{4},\mathcal{C}^{A}_{a},\mathcal{S}_{a}^{A})=3$; whereas, for $Bob$, $\mathcal{T}_{\varepsilon\iota}(t_{4},\mathcal{C}_{a},\mathcal{S}_{a}^{B})=7$
@@ 4.7 Utility function
For our scenario, upon the first negotiation iteration, $Bob$ is the sender, $Alice$ the receiver, and the first offer $o_{1}^{B\rightarrow A}=\langle\{t_{4}\},\{\langle elderly,$ $true\rangle\}\rangle$ is compared with the $Alice$’s $\bar{o}=\langle\{t_{4}\},$ $\{\langle injured,true\rangle,$ $\langle flight\_anxiety,true\rangle\}\rangle$ (see the $Alice$’s profile in the context $\mathcal{C}^{A}_{a}$ in Table 1). As per the grades in the table, the utility function has value $\mathcal{U}(o_{1},\bar{o})$ == $\mathcal{O}_{\varepsilon\iota}(o_{1},\mathcal{C}^{A}_{a})-\mathcal{O}_{\varepsilon\iota}(\bar{o},\mathcal{C}^{A}_{a})$ == 24=22-4=-2. From the point of view of $Alice$, by considering the only status conditions disclosed by $Bob$ at the first iteration, a negative utility means that RobAssist $B$ performing the offered task $t_{4}$ for $Bob$ is less ethical than RobAssist $A$ performing $t_{4}$ for $Alice$.